Hi,
I really need some help on this motion problem. Could you give the working out in a table format as well as systems of equations? Thank you!
"Gertrude rows her boat 24 miles downstream in 4 hours. In order to make the return trip upstream in the same amount of time, the rate of the boat in still water was doubled. Find the rate of the current and the rate that Gertrude was rowing downstream."

Answer by Boreal(15235)

speed is s
current is c
downstream, s+c=6 mph (24/4)
upstream, 2s-c=6 mph, same amount of time, so average speed is the same
2s-c=s+c
s=2c
2c+c=6, in first equation
c=2 mph ANSWER
so s=4 mph. ANSWER
6 mph downstream
2(4) or 8 mph upstream-2 mph current is 6 mph
For table speed* time= distance, and that may be filled in.